On July 1, 2016, Farm Fresh Industries purchased a specialized delivery truck for $205,800. At the time, Farm Fresh estimated the truck to have a useful life of eight years and a residual value of $33,000. On March 1, 2021, the truck was sold for $102,000. Farm Fresh uses the straight-line depreciation method for all of its plant and equipment. Partial-year depreciation is calculated based on the number of months the asset is in service.

Required:
1. Prepare the journal entry to update depreciation in 2021.
2. Prepare the journal entry to record the sale of the truck.
3. Assuming that the truck was instead sold for $118,000, prepare the journal entry to record the sale.
Answer #1

Accumulated depreciation = (205800-33000/8*4.5) = 97200

Jan 1 to Feb 28 depreciation = (205800-97200-33000)/3.5*2/12 = 3600

a) Journal entry

Date account and explanation Debit Credit
Mar 1 Depreciation expense 3600
Accumulated depreciation-Delivery truck 3600

b) Journal entry

Date account and explanation Debit Credit
Mar 1 cash 102000
Accumulated depreciation-delivery truck (97200+3600) 100800
Loss on sale of equipment 3000
Delivery truck 205800

c) Journal entry

Date account and explanation Debit Credit
Mar 1 cash 118000
Accumulated depreciation-delivery truck (97200+3600) 100800
gain on sale of equipment 13000
Delivery truck 205800